if w=9-4i then: 

if w = 9 - 4i then: 
                                                
a. Find the complex conjugate of w, denoted by w

b. Find 

 w + w
———————
   2

Found 2 solutions by ewatrrr, Edwin McCravy:
Answer by ewatrrr(24785)

 
Hi


b. Find (w+w)/2

18/2 = 9

Answer by Edwin McCravy(20063)


If w = 9 - 4i then: 
                                                
a. Find the complex conjugate of w, denoted by w  

Just change the sign of the term with "i".  So

w = 9 + 4i



b. Find 

 w + w
———————
   2

Substituting:

 (9 + 4i) + (9 - 4i)
—————————————————————  =
         2


      9 + 4i + 9 - 4i
     ————————————————— =
             2


                   18   
                  ———— =
                    2


                    9

Edwin